How much water must be vaporized from 0.4 molar 200mL H2SO4 solution to make it 1.6 molar at same temperature?
1
Expert's answer
2018-12-07T01:56:08-0500

Answer on Question #83609

How much water must be vaporized from 0.4 molar 200mL H2SO4 solution to make it 1.6 molar at same temperature?

Solution

Find n(H2SO4) in the first solution: n= c*V = 0.4 mol/L *0.2 L = 0.08 mol

In the second solution n(H2SO4) is the sane as in the first one, n(H2SO4) = 0.08 mol, but the volume of the solution desreases as water vaporizes: V(solution ) = 0.2 L - V(H2O).

In the second solution concentartion of H2SO4 is:

1.6 = 0.08 /(0.2 - V(H2O))

0.05 = 0.2 - V(H2O)

V(H2O) = 0.15 L =150 mL

Answer: 150 mL